OpenStudy (anonymous):

3/45 how to come up with answer

OpenStudy (anonymous):

what are you suppose to do? unless u mean \[3\div 45 =1/15 = .066666666666666666666666666666666\] and the 5 just continues infinitely

OpenStudy (anonymous):

3/45 = 1/15 If its division you type it in 3 / 45 = in calculator.